OpenStudy (anonymous):

what is the equation of the line that is parallel to the line 3x+4y+6=0 going through the point (2,2)?

OpenStudy (anonymous):

go get em

OpenStudy (anonymous):

btw did you see that i am now over 1000? maybe i will catch up to amistre?

myininaya (myininaya):

3x+4y+6=0 4y=-3x-6 y=-3/4*x-6/4 y=-3/4*x-3/2 slope of this line is -3/4 so slope of any parallel line is also -3/4 so line has form y=mx+b we know a ponint on this line (x=2,y=2) 2=-3/4*2+b solve for b 2+3/4*2=b so the equation of the line is y=-3/4*x+2+3/2

myininaya (myininaya):

yes satellite you will be the greatest